A consultant has collected the following information regarding Middle Road Publishing:

Total assets

$3,000 million

    Tax rate

40%

Operating income (EBIT)

$850 million

    Debt ratio

0%

Interest expense

$0 million

    WACC

10%

Net income

$510 million

    M/B ratio

1.00´

Share price

$32.00

   EPS = DPS

$3.20


The company has no growth opportunities (g = 0), so the company pays out all of its earnings as dividends (EPS = DPS). The consultant believes that if the company moves to a capital structure financed with 20 percent debt and 80 percent equity (based on market values) that the cost of equity will increase to 11 percent and that the pre-tax cost of debt will be 10 percent. If the company makes this change, what would be the total market value of the firm? (The answers are in millions.)

Select one:

a. $5,100

b. $4,600

c. $4,000

d. $3,200

e. $3,800
Answer #1

Debt Equity ratio :1:4 ,Cost of debt 10%

Solution Calculation of new WACC:

Capital Structure   capital    Proportion cost of capital WACC

Equity 2400 0.8 11   8.8

Debt 600     0.2 10(1-0.4) 1.2   

Total 3000 11  

Ke= 11

PE= MPS/EPS

=32/3.2

=10 times

No of Share Outstanding:

=(Net Income)/EPS

=($510)/3.2

= 159.375 Million

New Income Calculation : ('$ Million')

EBIT 850

Interest 60 (600*10%)

EBT 790

Tax @ 40% 316

PAT 474   

No of Shares o/s 159.375

New EPS 2.97

New Market Price Share: New EPS * PE

= 2.97*10 times

= 29.7